The complement of an angle measures 31°. What is the measure of the angle?
Answer by RAY100(1637)

Let x= angle,,,,,,,,,y= complement
.
x+y =90,,,,def complement
.
but y=31 degrees,,,,,given
.
x+31 = 90
.
x=90-31 = 59 degrees
.
.
check,,,,,, 31 +59 =90....ok
.